There is an ongoing incident on CP02, that our team currently investigating with the NOC team in Data Center.
UPDATE: There was a hardware failure on CP02 node earlier in the day. Our on site DC engineers have replaced node with a healthy one. All services are back online and we are monitoring
the new node. Happy Easter time.
The current server will again be replaced within the next few days with a complete new one as it is for emergencies purposes. You might feel your box run a bit less performant until the server is replaced with the new one.
Down Time: 03 hours and 20 minutes
Previous Uptime: 137 Days